OpenStudy (anonymous):

a baker used cups of batter to make muffins. it took 2/3 cup of batter to make 1 muffin. how many muffins did the baker make? please show how to set the problem up.

OpenStudy (anonymous):

Hi, \(\huge\color{lightblue}{{Welcome~To~Open~Study}!!!!!}\)

OpenStudy (anonymous):

Do you have any ideas?

OpenStudy (awesomeadam65):

\[\frac{ 1 }{ 1 } \times \frac{ 2 }{ 3 } =\]

OpenStudy (awesomeadam65):

or

OpenStudy (awesomeadam65):

\[\frac{ 3 }{ 2 } \times \frac{ 1 }{ 1 }\]